OpenStudy (anonymous):

x^-2/3 + 3x^-1/3 + 2 = 0

OpenStudy (anonymous):

parentheses would be appreciated

OpenStudy (anonymous):

(x ^{-2/3}) + (3x ^{-1/3}) + 2 =0

OpenStudy (anonymous):

lets use a different variable yeah

OpenStudy (anonymous):

let u=x^(-1/3) and the that means that x^(-2/3)=u^2

OpenStudy (anonymous):

can you solve it from there?

OpenStudy (anonymous):

I will give it a shot. Thank you for your help

OpenStudy (anonymous):

let me know what you get, or if you need more help

OpenStudy (anonymous):

i get u = -1, -2. When I plug them back into the equation -1 works, -2 does not the answer is supposed to be -1, -1/8. I don't understand the -1/8

OpenStudy (anonymous):

let me actually do the arithmetic, give me a sec

OpenStudy (anonymous):

sorry i took so long my friend. So when you solve that quadratic you get -1 and -2. Now we have to use the orginal substituion of u=x^(-1/3). This is what you do: u=x^(-1/3) x=u^-3 Now we plug in our value of -1 and -2 into u^-3 x=(-1)^-3 or x=(-2)^-3 x=-1 or x=-1/8

OpenStudy (anonymous):

thank you very very much

OpenStudy (anonymous):

no problem